Check the convergence of the sequence defined by 𝑒𝑛+1 = π‘Ž/ 1+𝑒𝑛 where π‘Ž > 0, 𝑒1 > 0.


1
Expert's answer
2022-01-03T09:07:11-0500

"u_{n+1}=1+\frac{1}{u}"


mapΒ "u\to 1+\frac{1}{u}"Β can be extended to a Moebius transformation of the Riemann sphere

"C\cup \{\infin\}:"

"z\to \frac{z+1}{z},T(0)=\infin,T(\infin)=1"

Its fixed points are:

"a=\frac{(1+\sqrt 5)}{2},b=\frac{(1-\sqrt 5)}{2}"

obtained by solving the equation

"z^2-z-1=0"

We now introduce a new complex coordinate w on C, related to z via

"w=\phi(z)=\frac{z-a}{z-b}\implies z=\phi^{-1}(w)=\frac{a-bw}{1-w}"

The fixed points now are w = 0 andΒ "w=\infin"

in terms of the new coordinate w the transformation T appears as

"\tilde{T}=\phi \circ T \circ \phi^{-1}"Β , then:

"\tilde{T}: w\to \frac{b}{a}w,\tilde{T}(0)=0,\tilde{T}(\infin)=\infin"


since

"\frac{b}{a}=\frac{3-\sqrt 5}{2}=-0.382"

we can infer that the fixed point 0 is attracting with basin of attraction all of C, whileΒ "\infin"Β is repelling. This allows to conclude that in the original setting all initial pointsΒ "u_0\neq b"

lead toΒ "\displaystyle \lim_{n\to \infin} u_n=a"


So, the sequence converges.